OpenStudy (anonymous):

Algebra help? 1.)Find the lowest common denominator. 1 over (x+2)^2 1 over (x-2)^2 2 over x^2-4 2.)Solve the following equation: 2/3+1/x=2/5-3/x 3.)Simplify this expression. 2 over x-y + 3 over x+y - 5 over y-x - 7x-9y over y^2 - x^2 4.)Simplify. 1-x/y over 1-x^2/y^2 5.) Solve the following word problem. One pipe can fill a tank in 36 minutes, a second can fill it in 9 minutes, and a third can fill it in 12 minutes. If the tank is empty, how long will the three pipes, operating together, take to fill it?

OpenStudy (anonymous):

6.)Solve the following equation: 1/3(y-2)-5/6(y+1)=3/4(y-3)-2 7.)Solve the following equation: 3x+1 over 3x+2 = x+2 over x-5 8.)Simplify. x+y - x^2-y^2 over x-y 9.)Solve the following word problem. A man travels from Town X to Town Y at an average rate of 60 mph and returns at an average rate of 50 mph. He takes a 1/2 hour longer than he would take if he made the round trip at an average of 55 mph. What is the distance from Town X to Town Y? 10.)Simplify: x/2+x/5+x/3 11.)Select an equivalent form of this equation: x/12 - 7=x/4 12.) Solve the following equation: (2 over 3-x)+(3 over x+3)= 6x over x^2-9